OpenStudy (anonymous):

What common factor should be written in the blank? 12y + 27 = ___(4y + 9)

OpenStudy (goldphenoix):

Well. What number multiply by 4y would give us 12y?

OpenStudy (anonymous):

3y?

OpenStudy (goldphenoix):

Not quite. 4y * 3 y = 12y^2. Since 4 * 3 = 12. And y * y= y^2.

OpenStudy (goldphenoix):

Alright. What number multiply by 9 would give us 27?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

would it be 9y/

OpenStudy (goldphenoix):

Hmm? Your saying 9y is a common factor?

OpenStudy (anonymous):

Im so confused...

OpenStudy (goldphenoix):

If so, then 4y * 9y = 36y^2. Before, you said that 3 * 9 =27. So do you think that 3 * 4y = 12y?

OpenStudy (anonymous):

yes

OpenStudy (goldphenoix):

Great. Therfore, 3 is a common factor. :)

OpenStudy (anonymous):

Thank you!!

OpenStudy (goldphenoix):

You can check your work if you want. 3(4y + 9) 3 * 4y = 12y. 3 * 9 =27. 12y + 27 = 12y + 27. I hope this help. :)